Grade 12th passMechanics

A girl is sitting with her dog at the left end of a boat of length 10.0 m. The mass of the
girl, her dog and the boat are 60.0 kg, 30.0 kg and 100.0 kg respectively. The boat is at
rest in the middle of the lake. Calculate the centre of mass of the system. If the dog
moves to the other end of the boat, the girl staying at the same place, how far and in
what direction does the boat move?

ROHIT SHARMA
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
If point X which measured from the left end of the boat is the centre of mass of the system:
x * (60 +30) = 100* ((10/2) - x)
90 x = 500 - 100 x
x = 500/190 = 2.63
2.63 m from the left end of the boat.
(b) 160 * Vb - 30* (Vd - Vb) = 0
 Vb /Vd = 3/19
 S = 10 * Vb/Vd = 1.57m
The boat is moving in the opposite direction to the dog’s movement 
Answer: The boat is moving in the opposite direction to the dog’s movement; the length which 
boat passes is 1.58m